For a constant cost industry in which all firms the same cost functions, their long-run average...

  1. For a constant cost industry in which all firms the same cost functions, their long-run average cost is minimized at $10 per unit output and 20 units (i.e. q = 20). Market demand is given by QD=DP=1,500-50PnT6uCt27h3TZsAXLf7boQoMRfl213ay6IwD8FLxu.
    1. Find the long-run market supply function
    2. Find the long-run equilibrium price (P*), market quantity (Q*), firm output (q*), number of firms (n), and each firm’s profit.
    3. The short-run total cost function associated with each firm’s long-run costs is SCq=0.5q2-10q+200UulmjTO8AAAAASUVORK5CYII=. Calculate the short-run average cost and short-run marginal cost functions. At what level of output is the short-run average cost function minimized?
    4. Calculate the short-run supply function for each firm and the market supply function. Assume the market is currently in the long run equilibrium from part (b)
    5. Suppose market demand shifts upward to QD=DP=2,000-50P6EGn6OadbN9C0COmVJemthaDXgMIZejdOSNzA6Ro. Find the equilibrium price, market quantity, and number of firms for the very short run when firms cannot change output.
    6. For the new market demand, find the short-run equilibrium price, market quantity, firm output, and firm level of profit.
    7. For the new market demand, find the long-run equilibrium price, market quantity, and number of firms.
Answer #1

mini AC = 10 at 12o-al. • Oo=150o-sop as long een supply is perfectly elastic (P = 10 In cagenem P=10, 9-20 do= 1500-50/200-5